英語 (English)

主題別英語 (Thematic English) (夜間主 ((化生知)1 年))

ワイスドナルド · 非常勤講師 / 全学共通教育センター 1 単位後期火 11 ・ 12

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 TESOL Goal 2, Standard 2 Students will use English to obtain, process, construct, and provide subject matter information in spoken and written form; listening to, speaking, reading, and writing about subject matter information; gathering information orally and in writing; retelling information; formulating and asking questions.

【授業の概要】 Topics: Each week we will read, talk, and write about one or two

of the biggest news stories in the world. 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. Read and understand news stories in English.

2. Write about what you have read and talk about it in class. 【授業の計画】

1. Introduction to the course; examples of news stories and how to write and talk about them.

2. News story of the week.

3. News story of the week.

4. News story of the week.

5. News story of the week.

6. News story of the week.

7. News story of the week.

8. News story of the week.

9. News story of the week.

10. News story of the week.

11. News story of the week.

12. News story of the week.

13. News story of the week.

14. Preparation for the final exam.

15. Final exam 16. Course review. 【教科書】 None. All material for the class will be taken from the Internet, from

English web sites. 【参考書等】 [参考資料] 【成績評価の方法】 Final exam 50% Homework assignments 25% Class participation

25%. 【再試験の有無】 [再評価] 【受講者へのメッセージ】 There will be reading, writing, listening, and speaking practice each

week. All homework will be sent in by email.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221371 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ワイス . (オフィスアワー: woosongdonald@yahoo.com)

English

Thematic English (夜間主 ((化生知)1 年))

Donald Weiss · Part-time Lecturer / Center for General Education 1 unit 後期火 11 ・ 12

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ TESOL Goal 2, Standard 2 Students will use English to obtain, process, construct, and provide subject matter information in spoken and written form; listening to, speaking, reading, and writing about subject matter information; gathering information orally and in writing; retelling information; formulating and asking questions.

Outline⟩ Topics: Each week we will read, talk, and write about one or two of the

biggest news stories in the world. �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. Read and understand news stories in English.

2. Write about what you have read and talk about it in class.

Schedule⟩

1. Introduction to the course; examples of news stories and how to write and talk about them.

2. News story of the week.

3. News story of the week.

4. News story of the week.

5. News story of the week.

6. News story of the week.

7. News story of the week.

8. News story of the week.

9. News story of the week.

10. News story of the week.

11. News story of the week.

12. News story of the week.

13. News story of the week.

14. Preparation for the final exam.

15. Final exam 16. Course review. Textbook⟩ None. All material for the class will be taken from the Internet, from

English web sites. Reference⟩ [参考資料] Evaluation Criteria⟩ Final exam 50% Homework assignments 25% Class

participation 25%. Re-evaluation⟩ [再評価] Message⟩ There will be reading, writing, listening, and speaking practice each

week. All homework will be sent in by email.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221371 Contact (Office-Hour, Room, E-mail)⟩

⇒ Weiss . (Office Hour: woosongdonald@yahoo.com)